OpenStudy (anonymous):

How many grams of glucose is in 25.0 mL of a 20% glucose solution mixed with 75.0 mL of water? Would it be 2.5 grams?

OpenStudy (anonymous):

20% = 1/5 so 25.0/5 = 5 then 5/2 = 2.5??